Please help me understand this scenario:

1. How can Bank X reduce its WACC?

2. How do I solve for EBT (Earnings before taxes) for the following hypothetical financial firm given the following information:

Tax rate = 25%

Net income = $12 million.

Also, if EBIT (Earnings before interest and taxes) = $21 million, what would the interest expense equal?
Answer #1

(2) Tax Rate = 25 %,Net Income = $ 12 million,

EBT (Earnings Before Tax) = Net Income / (1-Tax Rate) = 12 / (1-0.25) = $ 16 million

EBIT = EBT + Interest Expense

Interest Expense = EBIT - EBT = 21 - 12 = $ 9 million